Raymondʹs job pays $20 per hour for the first forty hours and $30 per hour for each hour in the
week worked above the 40 hours. If he earned $980 this week, how many overtime hours did he
work?

Answer by ikleyn(52781)

.

    40*20 + 30*x = 980

            30x  = 980 - 40*20

            30x  =    180

              x  =    180/30 = 6  hours overtime.    ANSWER

All equations are self-explanatory.